# Robust Pronoun Use Fidelity with English LLMs: Are they Reasoning, Repeating, or Just Biased?

### v1

@@ 6 Injecting an Introductory Context
As Figure 3 shows, all models can use a simple introduction better than chance (up to 0.95 with MosaicBERT), but not as well as humans, who achieve perfect performance. We also see improvements with increasing model scale, with the exception of ALBERT-v2. Model accuracy on accusative pronouns is statistically significantly higher than both possessive and nominative pronouns, but patterns across models and scaling are the same.
(a) With 0 distractors
(b) With 5 distractors
Figure 5: Scaling behaviour by architecture. With 0 distractors (above), encoder-only models are comparable to decoder-only models that are orders of magnitude larger. With 5 distractors (below), encoder-only models are far better than all decoder-only models.
Which pronouns are harder? Even in the simplest case of the pronoun use fidelity task, patterns emerge when split by pronoun, as shown in Figure 4. Overall model accuracy on he/him/his is significantly higher than she/her/her, which in turn is significantly higher than both they/them/their and xe/xem/xyr, in line with previous findings that language technology has gaps when it comes to gender-neutral pronouns, especially neopronouns (Lauscher et al., 2023). Models show intriguing patterns with these last two pronoun sets. Most encoder-only models appear to handle the neopronoun better than singular they (e.g., BERT-large has an accuracy of 0.78 on xe/xem/xyr compared to 0.60 on they/them/their), which warrants further investigation. Decoder-only models smaller than 6.7B parameters struggle with the neopronoun, with every OPT and Pythia model smaller than 2.7B parameters performing below chance, and in some cases (e.g., Pythia-14M, Pythia-70M and Pythia-160M) even performing close to 0.0. Beyond this scale, however, models perform better on xe/xem/xyr than on singular they, with Llama-13B achieving 0.96 accuracy on the neopronoun. These differences are statistically significant. As the training data for individual model families is the same, this might suggest that decoder-only models could learn to generalize to novel pronouns starting at the scale of 6.7B parameters. This could also explain the poor performance that some previous studies of neopronouns find, as the largest model that Hossain et al. (2023) experiments with, for instance, is OPT-6.7B. The lower performance of bigger models with singular they could also be a reflection of human processing difficulties with definite, specific (singular) they, as has been observed in linguistics (Conrod, 2019).
7 Adding Distractors
To further probe whether models actually use “reasoning” when supplied with an introductory context, we systematically inject sentences containing distractor pronouns between the introduction and the task, to examine how this affects performance.
[…]
Correct answer: her
[…]
Figure 5 compares results on the easiest case of pronoun use fidelity (no distractors) with the hardest case (5 distractors), split by architecture. Surprisingly, with no distractors, encoder-only models are much better than decoder-only models of the same scale, and their performance is comparable to or better than decoder-only models that are orders of magnitude larger; RoBERTa-base (125M) is 0.86 accurate compared to OPT-125M’s 0.55, and exceeds OPT-66B’s 0.83 despite being more than 500 times smaller. In the hardest version of our task with five distractors, encoder-only models are far better than all decoder-only models, which show dramatically degraded performance. Llama-70B only achieves 0.37 accuracy, compared to MosaicBERT’s impressive 0.87. The lack of robustness of decoder-only models to distractors is striking, given that most state-of-the-art models in popular use today are decoder-only models.
@@ 7 Adding Distractors
Figure 6: Model behaviour with additional distractors. Decoder-only models (above) get steadily worse, whereas encoder-only models (below) get worse with one distractor and then slowly recover, plateauing below their no-distractor performance.
Examining these results granularly, we see more interesting patterns across architectures and scales, shown in Figure 6. All decoder-only models get steadily worse as distractors are added, whereas encoder-only models perform the worst with one distractor and then seem to slowly recover, never quite reaching their level of performance with no distractors. Scaling seems to hold within model families, with larger models performing better with more distractors than smaller models of the same type.
Our results highlight the need to investigate task performance on a variety of architectures, as they might not pattern similarly. We hypothesize that the vastly different behaviour of encoder-only and decoder-only models with additional distractors reflects their different pre-training goals: encoder-only models are trained to learn rich representations of text, whereas decoder-only models are trained to predict the next token, which might make them more prone to recency bias.
Which cases and pronouns are harder with distractor sentences? With five distractor sentences, model accuracy with different pronoun sets follows the same statistically significant patterns as before. Interestingly, accusative pronoun performance is significantly different from the other cases as before, but in the opposite direction, i.e., worse than performance on nominative and possessive pronouns. This change from the best-performing pronoun case to the worst could be explained by a tendency for models to simply repeat the most recent accusative pronoun, a hypothesis we test next by examining model distractibility.
8 Distractibility versus Bias
In this section, we focus on the error cases for an in-depth analysis to confirm that models make mistakes because they are distracted by our distractor sentences. When a model gets the answer wrong, it is for one of three reasons: (1) distractibility, i.e., repeating the distractor pronoun, (2) bias, i.e., reverting to the model’s context-free prediction, or (3) picking a completely different pronoun. All three possibilities are illustrated below, and we hypothesize that the first two possibilities are much more frequent than the third.
In cases where the distractor pronoun is the same as the model’s context-free prediction, it is impossible to disentangle distractibility and bias just from the model’s prediction. Hence, we exclude these and focus on the unambiguous error cases. We find that 74-93% of unambiguous model errors can be attributed to the interplay of model distractibility and bias, confirming our hypothesis.
Refer to captionRefer to caption
@@ 8 Distractibility versus Bias
With introduction and distractors
@@
Are errors with distractor sentences actually due to distraction? We first examine model distractibility, i.e., what percentage of errors are caused by models repeating the distractor pronoun instead of the correct pronoun. As expected, Figure 7 shows that across models, distraction is indeed the primary type of error for most models. Decoder-only models get increasingly distractible with more distractors, i.e., the proportion of errors due to distractor pronoun repetition steadily increases as distractors are added, saturating just below 85%. On the other hand, encoder-only models seem to become less distractible with the addition of more distractors. We know from the previous section that encoder-only models recover in their pronoun use fidelity with 2-5 distractors, but here we measure distractibility as a percentage of all errors. Thus, a constant or increasing proportion of all the model errors could be due to distraction, and the fact that it isn’t for encoder-only models is quite surprising! We leave it to future work to investigate whether this behaviour is related to positional bias or context length.
When examining distractibility by pronoun case, we find that model distractibility is consistently higher with accusative pronouns compared to other cases. In combination, the high model performance on accusative pronouns with no distractors and the low performance with distractors strongly suggests that models tend to simply repeat the most recent pronoun, rather than reverting to biased predictions or using anything resembling robust “reasoning.”
As encoder-only models make proportionally fewer distraction errors, their errors due to bias go up. With BERT-large in particular, as soon as there is more than one distractor, the biggest proportion of errors is due to bias rather than distraction. BERT-large appears more biased and less distractible than BERT-base, in contrast to all other models. Generally, larger models seem to be more distractible and revert to their bias less often, whereas smaller models are more biased and less distractible. Our findings on bias errors contrast with Tal et al. (2022), where larger models make a higher proportion of bias errors on a downstream task than smaller models. We hypothesize the differences in our results could be explained by the presence of distractors in our task, which clearly have a very strong influence on model behaviour in this setting.
9 Prompting Shows Similar Results
Prompting is a different mechanism compared to evaluating log likelihoods and could in theory produce very different results. In practice, however, standard prompting has the same patterns of performance as before with an additional set of models, as shown in Figure 8.
(a) FLAN-T5
(b) Llama-2-chat
Figure 8: Performance of chat models with additional distractors, using prompting.
Bigger models are mostly better and degrade more gracefully than the smaller ones, but there remains a lot of variance across prompts, as shown in the box plots. We note that for Llama-2-chat, the best results with zero distractors are worse than the Llama-2 results with log likelihood evaluation. However, prompting seems to perform better with the addition of distractors, suggesting that chain-of-thought prompting—although excessive for a task this simple—might be a way of boosting performance.
@@ 10 Discussion and Future Work
Our results show that even the biggest models of today are not up to the task of pronoun use fidelity once it includes a single distractor sentence. All models are easily distracted, but encoder-only models and decoder-only models show very different patterns both in performance degradation with more distractors and their reasons for errors. Beyond architecture, model predictions across pronoun cases even in the absence of context exhibit a lot of variation, and performance on this type of reasoning task should be evaluated carefully, with attention to how the overall patterns break down by pronoun and case. Below we expand on some questions raised by our findings.
@@
“Reasoning” and “bias.” Throughout the paper, we refer to “reasoning” and “bias” but these are inaccurate terms, both as we use them and as NLP as a field does. What looks like “reasoning” when we inject an introductory context starts to look much more like repetition—or stochastic parroting (Bender et al., 2021)—when we add a distractor and see the same models performing drastically worse. Even the higher performance of encoder-only models cannot accurately be attributed to “reasoning” in the same way that we use this word for humans, as these models are not grounded in meaning from the real world (Bender and Koller, 2020). As these are all language models, it is more accurate to say that the way that decoder-only models model language is prone to repetition of recent examples of the same word class, compared to encoder-only models. As for “bias,” we show that statistical bias does not always overlap in predictable ways with social bias, by finding poor correlations across cases for model predictions. This has implications for NLP literature on model bias, where using sentence templates and nominative pronouns to quantify social bias in models is a popular methodology.
@@
Why exactly do we see the patterns we see? Our dataset design and error analysis show us how these models behave differently on the same task with a comparable evaluation method, and allow us to disentangle the effects of repetition, distraction and statistical bias. However, it is beyond the scope of this paper to investigate where in the model or data this comes from and what we can do about it. Tools from interpretability literature, e.g., attribution analysis, could help here, and are an important direction for future work.
What can we learn from humans? The author who validated the pronoun use fidelity task declared it easy and unambiguous, regardless of the number of distracting sentences. Clearly models do not correctly assign likelihood to the same extent, and prompting performance seems limited as well. One of the reasons for human success at this task might be our ability to assign pronouns to people and track them as separate individuals. Chain-of-thought prompting, or a different way of encouraging the model to explicitly track people and their pronouns, might therefore be a way of boosting model performance on this task.
@@
Beyond our dataset. Given the breadth of our task definition, future work could include examining pronoun use fidelity for participants, for names by extending Hossain et al. (2023), with differently ordered sentences, and in less artificial settings such as stories (Mostafazadeh et al., 2016) or with naturally collected data as in Webster et al. (2018) and Levy et al. (2021).
@@ 11 Related Work
The main areas of overlap with our work are: reasoning with pronouns, pronouns and occupational bias, and the effect of context on model predictions. The three studies that are closest to our work are Hossain et al. (2023) and Ovalle et al. (2023a), both of which explore pronoun reuse with an explicitly established coreference between an individual and a pronoun series, as well as Sharma et al. (2022), which injects context with an explicit coreference to encourage faithful pronoun translation. Unlike our work, however, none of these papers explore robustness via the addition of adversarial distractors.
Reasoning with pronouns. Most existing work about language model reasoning with pronouns focuses on the task of coreference resolution, typically using Winograd schemas (Levesque et al., 2012; Emelin and Sennrich, 2021). Of closest relevance to our work is Abdou et al. (2020), which shows that humans perform better than LLMs on perturbed versions of Winograd schemas designed to test generalization. Although Winograd schemas traditionally include reasoning about inanimate objects or groups, Winograd-like schemas that focus on third person singular pronouns and people entities have also been explored. The antecedents are sometimes named (Webster et al., 2018; Zhao et al., 2018) or referred to with a noun phrase, e.g., an occupation (Rudinger et al., 2018; Levy et al., 2021). Most studies focus on he and she, but recent work has expanded to include evaluation with singular they (Baumler and Rudinger, 2022) and with neopronouns (Cao and Daumé III, 2021; Felkner et al., 2023). Nevertheless, studies on coreference resolution evaluate whether systems can reason to identify the connection between a pronoun and an entity, without measuring whether LLMs can model faithful reuse of that pronoun later, as in our work.
Pronouns and occupational bias. Associations between pronouns and occupations have been studied extensively in NLP, in the subfields of fairness and machine translation. Many fairness studies examine intrinsic bias in masked token prediction (Kurita et al., 2019; de Vassimon Manela et al., 2021; Tal et al., 2022) in embeddings (Bolukbasi et al., 2016; Zhao et al., 2019; Gonen and Goldberg, 2019), and in translations (Stanovsky et al., 2019; Saunders et al., 2020; Iluz et al., 2023; Ghosh and Caliskan, 2023), particularly in context-aware machine translation (Müller et al., 2018; Voita et al., 2018; Fernandes et al., 2023). Work on translation bias with unambiguous sources and targets is closely connected to our English-only work, as there is a ground truth and therefore a way to measure accuracy and extrinsic bias.
The effect of context. The impact of context on the robustness of language model reasoning has been investigated in many areas other than pronoun fidelity, e.g., negation (Gubelmann and Handschuh, 2022), linguistic acceptability (Sinha et al., 2023), natural language inference (Srikanth and Rudinger, 2022), and question answering  (Liu et al., 2024; Levy et al., 2024).
@@ 12 Conclusion
We introduce the task of pronoun use fidelity and a corresponding dataset we designed to evaluate robust, faithful and harm-free pronoun use in language models, while disentangling the effects of distraction, repetition and bias. We find evidence of faithful pronoun reuse only in a very simple setting, where models are provided with an introductory context and no distractor pronouns. Even here, models show significant performance disparities with neopronouns, singular they and she/her/her, compared to he/him/his. Models are also not robust to more complex settings; even adding a single sentence containing a distractor pronoun, accuracy drops dramatically. Interestingly, encoder-only models are better at faithfully reusing pronouns than decoder-only models of the same scale, and they are also more robust to the addition of distractor sentences. We show that most model errors are indeed due to distraction, but with additional distractor sentences, encoder-only models become less distracted, while decoder-only models get even more distracted. Our results show that in a very simple setting for humans, widely-used large language models are unable to robustly and faithfully reason about pronouns, and continue to amplify discrimination against users of certain pronouns. We encourage researchers in bias and reasoning to bridge the performance gaps we report.
Limitations
Our task as it is defined in Section 2 is much broader than the scope of our dataset. We focus on occupations due to the wide attention they have received in prior literature, but we continue a long tradition of ignoring biases relating to the participants, e.g., child, taxpayer, etc. Additionally, we evaluate on a very easy version of this task that allows us to quantify repetition, i.e., the case of the pronoun required in the task is the same as the case shown in the context. Examining model performance where a pronoun is shown in one case and then elicited in a different case would be an interesting area for further investigation, particularly given our findings that models have uncorrelated statistical biases across different cases.
More broadly, much of the recent progress on reasoning datasets has been critically investigated and shown to often be a result of spurious correlations and dataset artifacts (Trichelair et al., 2019). Although the results in this paper show models struggling with our task, we caution readers that our dataset actually gives a very generous estimate of model reasoning performance for two reasons. Firstly, accuracy is a generous metric compared to consistency-based metrics across parallel versions of each data instance that only vary the pronoun that we consider (Elazar et al., 2021a, b). As we randomly sample from our generated data, we do not have exactly parallel versions to compute this metric. Secondly, and more importantly, many of our task sentences are not “Google-proof” (Levesque et al., 2012), i.e., they can be solved with shallow heuristics such as word co-occurrences. Consider the following task sentence: The janitor said not to step on the wet floor, otherwise ___ would have to mop it all over again. The association between janitor and mop is strong compared to child (the participant paired with janitor) and mop, and could easily be exploited by models to solve the dataset without solving the task with something resembling “reasoning.”
Another shallow heuristic that can be used to solve our current dataset is to simply return the first pronoun in the context, which happens to always be the correct answer. Our dataset design is flexible and allows for the creation of other orderings of sentences, but this is another example of why our dataset should only be used as an evaluation dataset, and models should not be pre-trained or fine-tuned with any splits of our data, nor provided with examples for in-context learning.
We take steps to prevent data contamination following Jacovi et al. (2023), and do not evaluate with models behind closed APIs that do not guarantee that our data will not be used to train future models. However, as we cannot guarantee an absence of data leakage unless we never release the dataset, we encourage caution in interpreting results on our dataset with models trained on data after March 2024.

### v2

@@ 6 Injecting an Introductory Context
As Figure 3 shows, all models perform better than chance at pronoun fidelity with a simple introduction (up to 0.95 with MosaicBERT), but not as well as humans, who achieve perfect performance. We also see improvements with increasing model scale, with the exception of ALBERT-v2, as in Tay et al. (2022).
Which pronouns are harder? Even in the simplest case of the pronoun fidelity task, patterns emerge when split by pronoun, as shown in Figure 4. Overall model accuracy on he/him/his is significantly higher than she/her/her, which in turn is significantly higher than both they/them/their and xe/xem/xyr, in line with previous findings that language technology has gaps when it comes to neopronouns (Lauscher et al., 2023). Models show intriguing patterns with these last two pronoun sets. Most encoder-only models appear to handle the neopronoun better than singular they (e.g., BERT-large has an accuracy of 0.78 on xe/xem/xyr compared to 0.60 on they/them/their), which warrants further investigation. Decoder-only models smaller than 6.7B parameters struggle with the neopronoun, with every OPT and Pythia model smaller than 2.7B parameters performing below chance, and in some cases (e.g., Pythia-14M, Pythia-70M and Pythia-160M) even performing close to 0.0. Beyond this scale, however, models perform better on xe/xem/xyr than on singular they, with Llama-13B achieving 0.96 accuracy on the neopronoun. These differences are statistically significant. As the training data for individual model families is the same, this might suggest that decoder-only models generalize to novel pronouns starting at the scale of 6.7B parameters, but in light of Schaeffer et al. (2023), this result could just as well be a mirage resulting from our use of accuracy, a discontinuous metric. In either case, our observations could also explain the poor performance that some previous studies of neopronouns find, as the largest model that Hossain et al. (2023) experiment with, for instance, is OPT-6.7B. The lower performance of bigger models with singular they could also be a reflection of human processing difficulties with definite, specific (singular) they, as has been observed in linguistics (Conrod, 2019).
@@
Figure 5: With more distractors, decoder-only models (above) get steadily worse; encoder-only models (below) get worse with one distractor and then recover, plateauing below their no-distractor accuracy.
(a) With 0 distractors
(b) With 5 distractors
Figure 6: Scaling behaviour by architecture. With 0 distractors (above), encoder-only models are comparable to decoder-only models orders of magnitude larger. With 5 distractors (below), encoder-only models are far better.
7 Adding Distractors
To further probe whether models actually “reason” when provided with context, we systematically inject sentences containing distractor pronouns between the introduction and the task, reflecting a natural usage scenario where multiple people are discussed with definite references and pronouns.
[…]
Correct answer: her
[…]
Figure 5 shows that distractors degrade performance for all models. Encoder-only and decoder-only models show different performance curves as more distractors are added: all decoder-only models get steadily worse, whereas encoder-only models perform the worst with one distractor and then seem to slowly recover, never quite reaching their level of performance with no distractors. Scaling generally holds within model families, with larger models performing better with more distractors than smaller models of the same type. Figure 6 examines the interplay of scaling and architecture at a higher level, comparing results on the easiest case of pronoun fidelity (no distractors) with the hardest case (5 distractors). Surprisingly, with no distractors, encoder-only models are much better than decoder-only models of the same scale, and their performance is comparable to or better than decoder-only models that are orders of magnitude larger; RoBERTa-base (125M) is 0.86 accurate compared to OPT-125M’s 0.55, and exceeds OPT-66B’s 0.83 despite being more than 500 times smaller. In the hardest version of our task with five distractors, encoder-only models are far better than all decoder-only models, which show dramatically degraded performance; Llama-70B only achieves 0.37 accuracy, compared to MosaicBERT’s impressive 0.87. The lack of robustness of decoder-only models to distractors is striking, given that most state-of-the-art models today are decoder-only models. We hypothesize that architectural differences might explain the performance gaps; encoder-only models might use bidirectional attention to more closely relate the entity mentions in the introduction and task sentences. Training on next token prediction might also make decoder-only models prone to recency bias.
Using vanilla and chain-of-thought prompting (Appendix I) show the same patterns of degradation, reinforcing that model pronoun fidelity is not robust, and good performance with no distractors (§6) is likely not due to “reasoning” at all.
@@ 7 Adding Distractors
8 Distractibility versus Bias
In adding distractor sentences, we add distance from the introduction via additional tokens that might make the model forget the original occupation-pronoun association, and the distractor pronoun also acts as a competing token that the model might accidentally repeat. In this section, we focus on the error cases to disentangle whether models are “forgetting” and reverting to biased predictions from Section 5, or if they are actually being distracted. When a model gets the answer wrong, it is for one of three reasons: (1) distractibility, i.e., repeating the distractor pronoun, (2) bias, i.e., reverting to the model’s context-free prediction, or (3) picking a completely different pronoun. All three possibilities are illustrated below , and we hypothesize that the first two possibilities are much more frequent than the third.
In cases where the distractor pronoun is the same as the model’s context-free prediction, it is impossible to disentangle distractibility and bias just from the model’s prediction. Hence, we exclude these and focus on the unambiguous error cases. As expected, we find that 74-93% of unambiguous model errors can be attributed to either model distractibility or bias.
@@ 8 Distractibility versus Bias
With introduction and distractors7)
@@
We first examine model distractibility, i.e., what percentage of errors are caused by models repeating the distractor pronoun instead of the correct pronoun. As expected, Figure 7 shows that across models, distraction is indeed the primary type of error for most models. Decoder-only models get increasingly distractible with more distractors, i.e., the proportion of errors due to distractor pronoun repetition steadily increases as distractors are added, saturating just below 85%. On the other hand, encoder-only models seem to become less distractible with the addition of more distractors. We know from the previous section that encoder-only models recover in their pronoun fidelity with 2-5 distractors, but here we measure distractibility as a percentage of all errors. Thus, a constant or increasing proportion of all the model errors could be due to distraction, and the fact that it isn’t for encoder-only models is quite surprising! We leave it to future work to investigate whether this behaviour relates to positional bias or context use.
As their proportion of distraction errors goes down, encoder-only models increasingly revert to biased predictions. With BERT-large in particular, as soon as there is more than one distractor, the biggest proportion of errors is due to bias rather than distraction. BERT-large appears more biased and less distractible than BERT-base, in contrast to all other models. Generally, larger models seem to be more distractible and revert to their bias less often, whereas smaller models are more biased and less distractible. Our findings on bias errors contrast with Tal et al. (2022), where larger models make a higher proportion of bias errors on a downstream task than smaller models. This might be due to our task having distractors, which seem to strongly influence model behaviour in this setting.
The high distractibility of all models strongly suggests that models simply repeat the most recent pronoun, rather than using anything resembling “reasoning,” and the contrast in error behaviour between encoder-only and decoder-only models further highlights their differences. This shows that claims about decoder-only models should not be applied to all LLMs, and that reasoning must be evaluated carefully, accounting for the possibility of inflated performance due to shallow heuristics like repetition.
@@ 9 Discussion and Future Work
Our results show that even the biggest models of today are not up to the task of pronoun fidelity once it includes a single sentence discussing another person. All models are easily distracted, but encoder-only models and decoder-only models show very different patterns both in performance degradation with more distractors and their reasons for errors. Performance on this type of reasoning task should be evaluated carefully, with attention to how the overall patterns break down by different pronouns and accounting for the possibility of repetition. Below we expand on some questions raised by our findings.
Improving robust pronoun fidelity. A natural direction of future work is to solve the problem of robust pronoun fidelity. We urge researchers interested in this direction to treat RUFF as an evaluation dataset, as it was designed. Due to the presence of positional and associative heuristics that we expand on in Appendix A, RUFF should not be seen as a source of data for fine-tuning or in-context learning, which is also why we do not run these experiments. Instead, a promising direction might be to encourage models to explicitly track associations between pronoun sets and individuals, just as people do.
@@
On “reasoning.” Throughout the paper, we refer to “reasoning,” but this is inaccurate. What looks like “reasoning” when we inject an introductory context starts to look much more like repetition—or stochastic parroting (Bender et al., 2021)—when we add a distractor and see the same models performing drastically worse. Even the higher performance of encoder-only models cannot accurately be attributed to “reasoning” in the same way that we use this word for humans, as these models are not grounded in meaning from the real world (Bender and Koller, 2020). We use the word reasoning in line with other work in the field, but note that as these are all language models, it is more accurate to say that the way that decoder-only models model language is prone to repetition of recent examples of the same word class, compared to encoder-only models.
@@
Why exactly do we see the patterns we see? Our dataset design and error analysis shed light on model behaviour, allowing us to evaluate different architectures comparably and disentangle the effects of repetition, distraction and statistical bias. However, it is beyond the scope of this paper to investigate where in the model architecture, neurons or pre-training data this comes from and what we can do about it towards improving reasoning and mitigating bias. Tools from interpretability literature, e.g., attribution analysis, could help here, and are an important direction for future work.
@@
Beyond our dataset. Given the breadth of our task definition, future work could include examining pronoun fidelity for participants, for names by extending Hossain et al. (2023), with differently ordered sentences, and in more natural settings such as stories (Mostafazadeh et al., 2016) or with real-world data as in Webster et al. (2018) and Levy et al. (2021). Additionally, we evaluate on a version of this task that allows us to quantify repetition, i.e., the grammatical case of the elicited pronoun is the same as the case shown in the context. Examining model performance where a pronoun is shown in one grammatical case and then elicited in a different one would be interesting to probe syntactic generalization.
@@ 10 Related Work
Pronoun fidelity.Hossain et al. (2023) and Ovalle et al. (2023a) both study pronoun fidelity when models are prompted with a pronoun series to use for an individual, but they only consider simplistic pronoun use with no more than one person at a time. Although we look at within-language pronoun use, faithful pronoun use in context has also been studied in machine translation (Müller et al., 2018; Voita et al., 2018; Fernandes et al., 2023), where there is also a ground truth. Similar to our work, Sharma et al. (2022) injects context with an explicit coreference to encourage faithful pronoun translation. However, none of these papers explore the robustness of pronoun fidelity in the presence of distractors.
Reasoning with pronouns. Most existing work about LLM reasoning with pronouns focuses on the task of coreference resolution, i.e., the ability to identify the connection between a pronoun and an entity, which may not translate to faithful reuse of that pronoun later, as in our work. Reasoning with pronouns typically uses Winograd schemas (Levesque et al., 2012; Abdou et al., 2020; Emelin and Sennrich, 2021), or Winograd-like schemas about named individuals (Webster et al., 2018; Zhao et al., 2018), or people referred to by their occupation (Rudinger et al., 2018; Levy et al., 2021). Most studies focus on he and she, but recent work has expanded to include singular they (Baumler and Rudinger, 2022) and with neopronouns (Cao and Daumé III, 2021; Felkner et al., 2023), as we do.
Pronouns and occupational bias. Stereotypical associations between pronouns and occupations have been studied in masked token prediction (Kurita et al., 2019; de Vassimon Manela et al., 2021; Tal et al., 2022) and embeddings (Bolukbasi et al., 2016; Zhao et al., 2019), but these studies typically use brittle methodology (Gonen and Goldberg, 2019; Seshadri et al., 2022) and measure intrinsic bias, which may not translate to extrinsic bias or harms (Goldfarb-Tarrant et al., 2021). Unlike these works, we evaluate extrinsic bias and performance through our focus on natural pronoun use in context.
Robustness in context. The impact of context on the robustness of language model reasoning has been investigated in many areas other than pronoun fidelity, e.g., negation (Gubelmann and Handschuh, 2022), linguistic acceptability (Sinha et al., 2023), natural language inference (Srikanth and Rudinger, 2022), and question answering  (Liu et al., 2024; Levy et al., 2024).
@@ 11 Conclusion
We introduce the task of pronoun fidelity to evaluate robust, faithful and harm-free pronoun use in language models, and we present RUFF, a dataset we designed to evaluate it. We find evidence of faithful pronoun use only in a very simple setting, i.e., where only one person is discussed. Even here, models show significant performance disparities with neopronouns, singular they and she/her/her, compared to he/him/his. Even adding a single sentence about a second individual with a different pronoun causes accuracy to drop dramatically, showing that pronoun fidelity is neither robust to non-adversarial distractors nor due to “reasoning.” As more distractor sentences are added, encoder-only models perform better overall, but increasingly revert to biased predictions, while decoder-only models get increasingly distracted. Our results show that in a setting that is very simple for humans, widely-used large language models are unable to robustly and faithfully reason about pronouns, and continue to amplify discrimination against users of certain pronouns. We encourage researchers to bridge the performance gaps we report and to more carefully evaluate “reasoning,” especially when simple repetition could inflate perceptions of model performance.